1. A ball with mass m kg is thrown upward with initial velocity 20 m/s from the roof of a building 23 m high. Neglect air resistance.

Use g = 9.8 m/s^2. Round your answers to one decimal place.

(a) Find the maximum height above the ground that the ball reaches.
x max = _______ meters (absolute tolerance +/- 0.1)

(b) Assuming that the ball misses the building on the way down, find the time that it hits the ground.

t tend =______ s (absolute tolerance +/- 0.1)

2. Determine whether the equation is exact. If it is exact, find the solution. If it is not exact, enter NS.

(3x+2)+(3y-3) y prime = 0
Do not enter an arbitary constant.
_________ = c, where c is an arbitrary constant.
1
Expert's answer
2020-06-17T18:55:00-0400

1(A) Height of the building is given 23m


Velocity with which ball is projected upward, "u = 20m\/s"

acceleration due to gravity is given, "g=9.8 m\/s^2"


At maximum height, velocity of ball will be zero. i.e. "v=0"

using equation, "v^2 = u^2 -2gh"


"0 = (20)^2 - 2*9.8*h"

"h = 20.4 m" above the building.

For height above the ground,

"H = 20.4 + 23 = 43.4m"


(B) time taken in upward motion, "t_u"

"v = u - gt_u"


putting values,

"0=20-9.8*t_u \implies t_u = \frac{20}{9.8}" s


Time taken in downward motion,

using equation, "H = \frac{1}{2} g t_d^2" as initial velocity for downward motion will be zero.

then

"t_d = \sqrt{\frac{2H}{g}} = \sqrt{\frac{2*43.4}{9.8}}"


total time for motion, "t= t_d + t_u = \frac{20}{9.8} + \sqrt{\frac{2*43.4}{9.8}}= 5.012 s"


2 Given equation is


"(3x+2) + (3y-3)\frac{dy}{dx} = 0"


"(3x+2){dx} + (3y-3){dy}= 0"

comparing equation with "Mdx + Ndy = 0"


"M = 3x+2" and "N = 3y-3"


For equation to be exact, it must follow,


"\frac{\partial M}{\partial y} = \frac{\partial N}{\partial x}"


"\frac{\partial M}{\partial y} = 0" and "\frac{\partial N}{\partial x} = 0"


Hence equation is exact.


Integrating both sides,


"\int (3x+2)dx + \int (3y-3)dy = C" where C is constant.


"\frac{3}{2} x^2 + 2x + \frac{3}{2} y^2 - 3y = C"